OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

First step: expand the brackets. 4 - x - 2 = 6 - 5x - 20 Second step: Collect like terms 2 - x = -5x - 14 Third step: Start rearranging to make x the subject. Let's start off with the real values then focus on x and its coefficients after. - x = -5x - 16 0 = -4x - 16 4x = -16 Can you work it out from here?